OpenStudy (anonymous):

A cargo plane flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220mph. The average speed on the way back was 200mph. How many hours did the trip there take?

OpenStudy (anonymous):

\[\frac {s}{t} = 220\] \[\frac{s}{t+1} = 200\] From the first equation: \[s = 220t\]From the second equation: \[s = 200t + 200\] Let's substitute s in the last equation with 220t and so we have: \[220t = 200t + 200\] From there, t = 10 which is the answer.